1) Jill purchased a share one year ago for $8.27, and it is now worth $14.48. The share paid a dividend of $1.20 during the year. What was the share's rate of return from capital appreciation during the year? (as a percentage to the nearest two decimal points. don't use % sign. eg 2.881% is 2.88)

2) You think that your chance of getting a well-paid job in an investment bank is about 5 per cent. If you get the job you will have a starting salary of $81,000 per year. However, if you don't make the cut then you will work at a fast food outlet for $40,000 per year. What is your expected starting salary? (to the nearest dollar)

Select one:

a. $42050

b. $78950

c. $81000

d. $40000
Answer #1

1 ) P1= $ 14.48

P0= $ 8.27

D= $ 1.2

Rate of return= (Total Benefit/ P0)*100

Total benefit= (P1-P0)+D

= $ 14.48- $ 8.27 + $ 1.2

= $ 7.41

Rate of Return = ($ 7.41/$ 8.27)*100

=89.60

2) Probabilty of getting well paid Job= 5% (P1)

Probabilty of not getting a well paid job = 1- P1

= 1 - 0.05 = 0.95 i.e 95% (P2)

Salary if you get the well paid Job = $ 81,000

Salary if you not get the well paid job = $ 40,000

Salary Probability Outcome

81,000

5% 81000*5%=4050
40,000 95% 40,000*95%=38000

Expecting Starting Salary = $ 4,050 + $ 38,000

= $ 42,050 /- (i.e:- a)